OpenStudy (anonymous):

Find the intercepts and then graph the line 6x+5y=-30

OpenStudy (anonymous):

This is what I got for the intercept x= (5,0) y= (0,6)

OpenStudy (anonymous):

x= (5,0) y= (0,-6) that was my final answe need help in graphing

OpenStudy (anonymous):

x-intercept: (−5,0) y-intercept: (0,−6)

OpenStudy (anonymous):

do u want me to explain?

OpenStudy (ankit042):

for find intercepts put x=0 solve for y and then put y=0 and solve for x Now if you will get two points, graph them and join then to get you line

OpenStudy (anonymous):

I got the intercept right just need in graphing it

OpenStudy (anonymous):

But I got this for x intercept (5,0)

OpenStudy (anonymous):

To find the x-intercept, substitute in 0 for y and solve for x. 6x+5(0)=−30 Multiply 5 by each term inside the parentheses. More Detail Multiply 5 by the 0 inside the parentheses. 6x+5⋅0=−30 Multiply 5 by 0 to get 0. 6x+0=−30 6x+0=−30 Remove the 0 from the polynomial; adding or subtracting 0 does not change the value of the expression. 6x=−30 Divide each term in the equation by 6. 6x/6=−30/6 Simplify the left-hand side of the equation by canceling the common factors. x=−30/6 Simplify the equation. Cancel the common factor of 6 in −30/6. Reduce the expression −30/6 by removing a factor of 6 from the numerator and denominator. x=−5 x=−5 To find the y-intercept, substitute in 0 for x and solve for y. 6(0)+5y=−30 Multiply 6 by each term inside the parentheses. 0+5y=−30 Remove the 0 from the polynomial; adding or subtracting 0 does not change the value of the expression. 5y=−30 Divide each term in the equation by 5. Simplify the left-hand side of the equation by canceling the common factors. Cancel the common factor of 5 in 5y/5. Reduce the expression 5y/5 by removing a factor of 5 from the numerator and denominator. y=−305 y=−30/5 Solve the equation. Cancel the common factor of 5 in −30/5. Reduce the expression −305 by removing a factor of 5 from the numerator and denominator. y=−6 y=−6 These are the x and y intercepts of the equation 6x+5y=−30. x-intercept: (−5,0) y-intercept: (0,−6)

OpenStudy (anonymous):

plug in my answers and see if their correct

OpenStudy (anonymous):

Ok I will plug in now thanks

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

Then the last one

OpenStudy (anonymous):

now u would be right if the equation was –6x + 5y = –30